# DynaSpec: Context-aware Dynamic Speculative Sampling for Large-Vocabulary Language Models

### v2

@@ 4.1 Preliminary
Target LLM. Let $\mathcal{T}$ represent the target LLM over vocabulary $\mathcal{V}$. Given an input sequence $x_{1:n}$, the embedding layer produces
@@
Applying $L$ transformer blocks $\mathcal{F}_{1},\ldots,\mathcal{F}_{L}$ yields
$H\;=\;(\mathcal{F}_{L}\circ\cdots\circ\mathcal{F}_{1})(Z_{0})\in\mathbb{R}^{n\times d}.$
An LM head with weights $W_{LM}\in\mathbb{R}^{d\times|\mathcal{V}|}$ maps to logits $O=HW_{LM}$, and the token probabilities are
$p_{\mathcal{T}}(y_{i}\mid x_{\leq i})\;=\;\operatorname{softmax}(O_{i}).$
Full Vocabulary Drafter. Speculative decoding methods (more specifically EAGLE (li2024eagle; li2024eagle2) series of models) train a lightweight draft model $\mathcal{D}$ (one transformer block, same latent width $d$) to approximate $\mathcal{T}$ at a lower cost. $\mathcal{D}$ reuses and freezes the embedding $\mathbf{E}$ and LM head $W_{LM}$ from $\mathcal{T}$; only its single transformer block is trained:
$\tilde{H}\;=\;\mathcal{D}\!\big(\mathbf{E}(x_{1:n})\big),\qquad\tilde{p}_{\text{EAGLE}}\;=\;\operatorname{softmax}(\tilde{H}W_{LM}).$
Static Sub-vocabulary Drafter. FR-Spec (zhao-etal-2025-fr) uses a frequency-ranked drafting with vocabulary $\mathcal{V}_{\text{high}}\!\subset\!\mathcal{V}$ the subset of high-frequency tokens (from corpus statistics). Concretely, select the corresponding LM-head columns.
$\widetilde{W}\in\mathbb{R}^{d\times|\mathcal{V}_{\text{high}}|},\quad\widetilde{W}[:,j]=W_{LM}[:,\,\mathcal{V}_{\text{high}}[j]],$
and compute
@@
4.2 Meta classifiers: Context-aware token Selection
We use a lightweight meta-classifier (interchangeably, the router) to select coarse token clusters (“meta-classes”) per context; details follow.
@@ 4.2 Meta classifiers: Context-aware token Selection
We partition the vocabulary $\mathcal{V}$ into $M$ coarse token clusters $\Pi\triangleq\{C_{1},\ldots,C_{M}\}$ via a clustering map $\tau:\mathcal{V}\!\to\![M]$. Following extreme-classification style routing, clusters are formed offline by spherical $k$-means on column-normalized LM-head weights $\{W_{LM}[:,v]/\|W_{LM}[:,v]\|_{2}\}_{v\in\mathcal{V}}$. We do not enforce balance (i.e., $|C_{m}|$ may vary), which empirically preserves recall for semantically dense regions.
@@
Given the token embedding $\mathbf{E}(x_{t})$ and the drafter’s previous hidden features $\tilde{H}_{t-1}$, a lightweight MLP $r_{\theta}:\mathbb{R}^{d_{r}}\!\to\!\mathbb{R}^{M}$ produces scores $s=r_{\theta}([\mathbf{E}(x_{t}),\tilde{H}_{t-1}])$ over clusters. The router runs on a separate parallel CUDA stream and completes while the drafter’s attention/MLP is executing, making its wall-clock overhead negligible.
Position-aware cluster selection and shortlist construction.
@@
$k_{\mathrm{c}}(t)=\begin{cases}k_{0},&t=0,\\[4.0pt] k_{1},&t=1,\\[4.0pt] \left\lfloor\dfrac{k_{1}+t}{2t}\right\rfloor,&t\geq 2,\end{cases}$
where $k_{0}>k_{1}$ and both are hyperparameters controlling the initial cluster budgets. Given router scores $s$, we select $\mathcal{K}(c,t)=\operatorname{TopK}_{k_{\mathrm{c}}(t)}(s)$ and construct the shortlist $\mathcal{V}_{S}(c,t)=\bigcup_{m\in\mathcal{K}(c,t)}C_{m}$.
@@
We cast routing as multi-label prediction over $M$ clusters with hard labels. Given the hidden state $\tilde{H}_{t}$ of the drafter, let $S^{+}$ denote the positive samples from the drafter model with full vocabulary at draft step $t$ (e.g., the top-$L$ tokens at step $t$). A cluster is positive iff it contains at least one positive token:
$y_{m}\;=\;\mathbb{I}\!\big[C_{m}\cap S^{+}\neq\varnothing\big]\in\{0,1\}.$
We train with binary cross-entropy over all meta labels where $\hat{y}_{m}=\sigma\!\big(s_{m})\big)$:
$\mathcal{L}_{\mathrm{r}}\;=\;\mathbb{E}\!\left[\sum_{m=1}^{M}\Big(-\,y_{m}\log\hat{y}_{m}-(1-y_{m})\log\big(1-\hat{y}_{m}\big)\Big)\right]$
Further training details are provided in A.2.
@@ 4.3 Dynamic Drafting
The drafting process is depicted in Figure 1. At each step $t$, the router (meta-classifier) selects the shortlist indices, while the drafter generates the hidden state for that step. A single gathered-head GEMM is then performed over the shortlisted tokens using this hidden state. The router runs on a separate CUDA stream and typically finishes before the drafter block as shown in Figure 2. The total draft time for a single pass can be expressed as,
@@
where $T_{\text{meta}}\!\ll\!T_{\text{core}}$ and $T_{\text{emb}}$ is generally small. Thus the only vocabulary-dependent term is the $d\times B(c,t)$ head compute (with $B(c,t)\!\ll\!|\mathcal{V}|$), reducing the denominator in Eq. 2 while preserving acceptance via context-aware support. Since $B(c,t)\!\propto\!|\mathcal{V}_{S}(c,t)|$ and the position-aware budget $k_{\mathrm{c}}(t)$ decay with $t$, the shortlist and thus the head compute shrink as generation proceeds, while larger early budgets preserve acceptance. In our implementation, index selection and the head GEMM are fused into a single custom CUDA kernel to avoid intermediate buffers and redundant global-memory traffic: the kernel performs warp-level column gathers with asynchronous, double-buffered shared-memory tiling, computes on Tensor Cores, and uses a persistent L2 access-policy window to keep activations cached. Verification is unchanged from standard speculative decoding.
@@ 5.1 Experimental Settings
We adhere to the experimental setup and evaluation of the EAGLE (li2024eagle; li2024eagle2) family and FR-Spec (zhao-etal-2025-fr). We implement the DynaSpec in the FR-Spec  [^2] https://github.com/thunlp/FR-Spec framework. All components are implemented in native C++ and CUDA with FR-Spec as the baseline, since Python call overhead would otherwise negate the latency gains from draft-time acceleration.
Datasets and Models. We benchmark DynaSpec over a diverse mix of seven tasks. Spec-Bench (xia2024unlocking) supplies six of them—Machine Translation (WMT14 DE to EN) (bojar2014findings), Multi-turn Conversation (MT-Bench) (zheng2023eric), RAG and QA from Natural Questions (kwiatkowski2019natural), Math from GSM8K (cobbe2021training), and Summarization (CNN/DailyMail) (nallapati2016abstractive)—each evaluated on 80 prompts. The seventh task is Code, for which we employ HumanEval (chen2021evaluating) with 164 problems. We bound generations to $1024$ tokens for Spec-Bench subtasks and HumanEval. We used Llama series of models, specifically Llama-33-8B-Instruct, which uses a vocabulary size of $128K$ for our experiments and Qwen-2-7B.
Evaluation metrics and Hardware Settings. We evaluate using metrics: #Mean Acceptance Length, the average number of tokens committed per draft–verification cycle. #Mean Acceptance Length is hardware-agnostic but does not capture the overhead of drafting. Our primary baseline is FR-Spec (zhao-etal-2025-fr), a static subset–selection approach built on EAGLE-2 (li2024eagle2); in contrast, DynaSpec performs dynamic subset selection atop EAGLE-2 (li2024eagle2). All experiments were run on a single NVIDIA A6000 GPU.
Model / Config MT Conv. RAG Math QA Summ. Code Average Llama-3-8B Full Vocab $3.65$ $4.13$ $4.04$ $4.32$ $3.46$ $3.69$ $3.92$ $3.89$ vocab size=128k FR-Spec $3.39$ $3.88$ $3.86$ $4.16$ $3.33$ $3.52$ $3.31$ $3.64$ vocab size=32k DynaSpec $3.51$ $4.07$ $3.93$ $4.23$ $3.38$ $3.55$ $3.85$ $3.79$ vocab size 27,477 26,989 27,358 27,210 27,019 26,933 26,168 27,022 Qwen-2-7B Full Vocab $2.92$ $3.79$ $3.43$ $4.34$ $3.07$ $3.43$ $4.17$ $3.59$ vocab size=152k FR-Spec $2.84$ $3.52$ $3.30$ $4.17$ $2.95$ $3.32$ $3.52$ $3.37$ vocab size=32k DynaSpec $2.86$ $3.72$ $3.32$ $4.18$ $2.97$ $3.24$ $3.96$ $3.46$ vocab size=27k
Table 1: Mean accepted tokens for Llama-3-8B and Qwen-2-7B across seven datasets. vocab size indicates the shortlisted number of tokens. Full Vocab corresponds to the EAGLE-2 baseline; FR-Spec uses 32k. For DynaSpec, shortlist size varies by dataset; per-dataset means are shown individually.
@@
Figure 3: Mean accepted length comparison for fixed top-$k$ clusters represented by DynaSpec-F vs. Position-aware top-$k$ clusters represented by DynaSpec-PA across seven datasets using Llama-3-8B model. Mean vocab size represents the mean shortlisted vocab size.

### v3

@@ 4.1 Notation and preliminaries
Target LLM. Let $T$ represent the target LLM over vocabulary of tokens’ vocabulary $V$. Given $d$-dimensional token embeddings $\mathbf{E}$, for an input sequence $x_{1:n}$, the embedding layer produces
@@
Applying $\ell$ transformer blocks $\mathcal{F}_{1},\ldots,\mathcal{F}_{\ell}$ yields
$H\;=\;(\mathcal{F}_{\ell}\circ\cdots\circ\mathcal{F}_{1})(Z_{0})\in\mathbb{R}^{n\times d}.$
An LM head with weights $W_{\mathrm{LM}}\in\mathbb{R}^{d\times|V|}$ maps to logits $O=HW_{\mathrm{LM}}$, and the token probabilities are
$p_{T}(y_{i}\mid x_{\leq i})\;=\;\operatorname{softmax}(O_{i}).$
Full Vocabulary Drafter. Speculative decoding methods (more specifically Eagle (Li et al., 2024b, a) series of models) train a lightweight draft model $D$ (one transformer block, same latent width $d$) to approximate $T$ at a lower cost. The drafter reuses and freezes the embedding $\mathbf{E}$ and LM head $W_{\mathrm{LM}}$ from $T$, and only its single transformer block is trained. The output distribution $q(.)$ over the tokens is computed as : $\tilde{H}\;=\;D\!\big(\mathbf{E}(x_{1:n})\big)$ and $q_{\text{Eagle}}\;=\;\operatorname{softmax}(\tilde{H}W_{\mathrm{LM}})$.
Static Sub-vocabulary Drafter. Fr-Spec (Zhao et al., 2025) uses a frequency-ranked drafting with vocabulary $V_{\text{high}}\!\subset\!V$ the subset of high-frequency tokens (from corpus statistics). Concretely, select the corresponding LM-head columns.
$\widetilde{W}\in\mathbb{R}^{d\times|V_{\text{high}}|},\quad\widetilde{W}[:,j]=W_{\mathrm{LM}}[:,\,V_{\text{high}}[j]],$
@@
4.2 Context-aware dynamic sub-vocabulary selection
@@
We partition the vocabulary $V$ into $M$ disjoint token clusters i.e. $V=\biguplus_{m=1}^{M}C_{m}$. These clusters are formed offline by spherical $k$-means on column-normalized LM-head weights which is given by :
$\{W_{LM}[:,v]/\|W_{LM}[:,v]\|_{2}\}_{v\in V}$. Since some tokens in the vocabulary share similar semantics, we can cluster them and use a meta-classifier to efficiently select a relevant subset. We do not enforce balance (i.e., $|C_{m}|$ may vary), which empirically preserves recall for semantically dense regions.
@@
Given the token embedding $\mathbf{E}(x_{t})$ and the drafter’s previous hidden features $\tilde{H}_{t-1}$, a lightweight MLP $r_{\theta}:\mathbb{R}^{d_{r}}\!\to\!\mathbb{R}^{M}$ produces scores $s=r_{\theta}([\mathbf{E}(x_{t}),\tilde{H}_{t-1}])$ over clusters. Given these scores, we select $\mathcal{K}(c,t)=\operatorname{TopK}_{k_{\mathrm{c}}(t)}(s)$ by a top-$k$ operation, and construct the shortlist $\mathcal{V}_{S}(c,t)=\biguplus_{m\in\mathcal{K}(c,t)}C_{m}$. At each draft step, $x_{t}$ may consist of a single token or multiple tokens, depending on the beam size at step $t$. When $x_{t}$ contains multiple tokens, we take the union of the clusters predicted for each token.
@@
$k_{\mathrm{c}}(t)=\begin{cases}k_{0},&t=0,\\[4.0pt] k_{1},&t=1,\\[4.0pt] \left\lfloor\dfrac{k_{1}+t}{2t}\right\rfloor,&t\geq 2,\end{cases}$ (3)
where $k_{0}>k_{1}$ and both are hyperparameters controlling the initial cluster budgets. As shown in Figure 1, the router runs on a separate parallel CUDA stream and completes while the drafter’s attention/MLP is executing, making its wall-clock overhead negligible. This is because the draft model is typically very lightweight and cannot fully utilize the GPU, allowing the meta-selection to run in parallel on a separate stream.
@@
We cast training the router $r_{\theta}$ as a multi-label classifier over $M$ clusters with hard labels. At each time step $t$, the router observes the current context features (we use the embedding of the current token and the previous drafter hidden state) and outputs a score vector :
$s_{t}=r_{\theta}([\textbf{E}(x_{t}),\tilde{H}_{t-1]}])\in\mathbb{R}^{M},\hskip 7.22743pt\hat{y}_{t,m}=\sigma(s_{t,m})\in(0,1)$
where $M$ is the number of clusters and $\sigma$ is the sigmoid. The router is implemented as a two-layer feed-forward neural network with ReLU activation. Given an input feature vector, the network projects it into a hidden representation and then maps it to cluster logits through a linear layer.
To construct the training labels, we run the Eagle-2 pipeline on a large set of prompts (e.g., ShareGPT and UltraChat), and at each step $t$, we record : (i) the drafter hidden state $\tilde{H}_{t-1}$, (ii) the current token embedding $\textbf{E}(x_{t})$, and (iii) a set of plausible next token $S_{t}^{+}\subset V$ (e.g., the top-$L$ tokens under the full-vocabulary drafter at that step). Each token $v\in V$ belongs to exactly one cluster $C_{m}$. We therefore induce a multi-hot target over clusters :
$y_{t,m}=\mathbb{I}[C_{m}\cap S_{t}^{+}\neq\emptyset]\in\{0,1\}.$
Intuitively, a cluster is labeled positive if it contains at least one high-probability token for that context. We optimize binary cross-entropy (independent logistic losses) over clusters:
$\mathcal{L}_{\mathrm{r}}\hskip-3.61371pt\;=\;\hskip-3.61371pt\mathbb{E}\!\left[\sum_{m=1}^{M}\hskip-3.61371pt\Big(\hskip-3.61371pt-\,y_{t,m}\log\hat{y}_{t,m}\hskip-3.61371pt-(1-y_{t,m})\log\big(1-\hat{y}_{t,m}\big)\Big)\right]$
It must be noted that even though the next token prediction is a multi-class problem, the Dynaspec’s router is trained as multi-label classifier. This is because at time step $t$, the supervision signal is a set of tokens $S_{t}^{+}$, which land in multiple clusters because they include synonyms, formatting variants, punctuation, code tokens, domain terms, etc. Therefore, the induced label set $Y_{t}=\{m:C_{m}\cap S_{t}^{+}\neq\emptyset\}$ usually has $|Y_{t}|>1$, making it a multi-label problem.
@@ 4.3 Dynamic Drafting
The drafting process for a single drafting step is depicted in Figure 2. Given a context $c$ at position $t$, the router first produces cluster scores and selects $\mathcal{K}(c,t)$, which induces the shortlist $\mathcal{V}_{S}(c,t)$. In parallel, the drafter transformer computes the step hidden state $\tilde{H}_{t}$. We execute routing on a separate CUDA stream so that router latency overlaps with the drafter’s attention/MLP computation; in practice $T_{meta}\ll T_{core}$ (c.f. 1), so routing is typically hidden by the draft forward pass.
After synchronizing the two streams, we compute drafter logits using a gathered LM head over the shortlisted tokens only. Let $B(c,t)=|\mathcal{V}_{S}(c,t)|$. Instead of the full projection $\tilde{H}_{t}W_{\text{LM}}$ with cost $\mathcal{O}(|V|d)$, we gather the $B(c,t)$ corresponding columns of $W_{\text{LM}}$ and perform a single GEMM of cost $\mathcal{O}(B(c,t)d)$ where $B(c,t)\ll|V|$. This yields the following per-step latency approximation:
@@
where $\max(.,.)$ term reflects the stream overlap between routing and the drafter core.
Since $T_{\text{index}+\text{gemm}}$ grows with $B(c,t)$, we use a position-aware cluster budget $k_{\mathrm{c}}(t)$ decay with $t$, allocating larger shortlists early (to protect acceptance) and smaller shortlists later (to reduce compute).
@@
In our implementation, index selection and the head GEMM are fused into a single custom CUDA kernel to avoid intermediate buffers and redundant global-memory traffic: the kernel performs warp-level column gathers with asynchronous, double-buffered shared-memory tiling, computes on Tensor Cores, and uses a persistent L2 access-policy window to keep activations cached. Verification is unchanged from standard speculative decoding.
In the next section, we present a general theoretical analysis of speculative decoding when the draft model is restricted to a vocabulary subset. Apart from the experimental comparisons on benchmark datasets later in the paper, the theoretical framework can also be used to analytically compare concrete variants such as DynaSpec and Fr-Spec.
5 Theoretical Analysis
For vocabulary $V$, at a decoding step with context $c$, let the target distribution be denoted as $p_{c}(x)\coloneqq p(x\mid c),x\in V$, and the full vocabulary draft distribution be denoted as $q_{c}(x)\coloneqq q(x\mid c)$. Note that this is only conceptual in our setting and corresponds to Eagle-style drafter if it had a full head. DynaSpec’s router chooses a context-dependent shortlist $S(c)\subset V$. The router selects a set of cluster indices $\mathcal{K}(c)\subset[M],|\mathcal{K}(c)|=k$, and tokens are shortlisted in as $S(c)=\cup_{m\in\mathcal{K}(c)}C_{m}$, also in context-dependent manner. The restricted and renormalized draft distribution on $S$ is given by (denoted as $q_{S,c}(x)$) :
$\frac{q_{c}(x)\mathbb{I}[x\in S(c)]}{q_{c}(S(c))}\eqqcolon q_{S,c}(x)$
Define a tail distribution $\omega_{c}(\cdot)$ supported on $V\setminus S(c)$. Then the actual proposal distribution is given by :
$\tilde{q}_{c}(x)\coloneqq(1-\beta)q_{S,c}(x)+\beta\omega_{c}(x)$
with $\beta\in(0,1)$ but with a small value. The tail-distribution $\omega_{c}(x)$ makes the proposal well-defined (full support) and lets us reason about correctness and acceptance. Also, it cleanly separates “shortlist quality” (through $q_{S}$) from the safety tail $(\beta,\omega)$, which is important for modelling purposes. Note that, $\omega_{c}(\cdot)$ can be a uniform distribution over the complement or a cheap approximate tail distribution. In practice, even if the implementation does not explicitly sample from a tail, we can set $\beta$ to an abstract “escape” probability that models either (a) fallbacks or (b) the effect of verifying and then resampling at rejection, making the theory robust to these details.
For one speculative step (one position), let $p_{c}(x)$ be the target next-token distribution given context $c$, and $\tilde{q}_{c}(x)$ be the effective draft proposal distribution for that step. The overlap between the two is denoted by $A(c)$ :
$A(c)\coloneqq\sum_{x\in V}\min(p_{c}(x),\tilde{q}_{c}(x))=1-\text{TV}(p_{c}(x),\tilde{q}_{c}(x))$ (4)
where $\text{TV}(.,.)$ is the total variation distance between distributions given by :
$\text{TV}(p,q)\coloneqq\frac{1}{2}\sum_{x\in V}|p(x)-q(x)|$ (5)
Write shorthand for $S(c)$ as : $S(c)=S$, and denoting by $p_{c}(S)\coloneqq\sum_{x\in S(c)}p_{c}(x)$, the target mass retained by the router shortlist $S$. Let $p_{c}(\cdot\mid S)$ and $q_{c}(\cdot\mid S)$ are the target and drafter distribution conditioned on the restriction to the shortlist $S$, i.e. $p_{c}(x\mid S)=p_{c}(x)/p_{c}(S)$ if $x\in S$, and 0 otherwise. Similarly, for $q_{c}(\cdot\mid S)$, start with the drafter’s full-vocabulary distribution $q_{c}(x),x\in V$(conceptual in our scenario), and compute its mass on $S$, i.e., $q_{c}(S)\coloneqq\sum_{x\in S(c)}q_{c}(x)$, then $q_{c}(x\mid S)=q_{c}(x)/q_{c}(S)$ if $x\in S$, and 0 otherwise.
Theorem 5.1 (Distribution overlap lower bound).
For any context $c$,
$A(c)\geq(1-\beta)\left(p_{c}(S)-\textnormal{TV}(p_{c}(\cdot\mid S),q_{c}(\cdot\mid S))\right).$
The proof is given appendix A.2. Essentially, the theorem above suggests that two issues can hurt acceptance. Firstly, coverage problem which is captured by $p_{c}(S)$, it is small when most target mass lies outside $S$. Secondly, even if $p_{c}(S)$ is large, the drafter might put probability on the wrong token within $S$, which is the quantity captured by $\text{TV}(p_{c}(\cdot\mid S),q_{c}(\cdot\mid S))$. Concretely, using the corresponding shortlists computed by Dynaspec and FR-Spec, i.e., $S_{\text{Dyn}}$ and $S_{\text{Fr}}$ respectively, one can compute the key quantities $p_{c}(S)$ and $\textnormal{TV}(p_{c}(\cdot\mid S),q_{c}(\cdot\mid S))$ in the above theorem 5.1 and compare the overlap lower bounds. These quantities are computed for the HumanEval (Code) and compared for the Dynaspec and FR-Spec in Table 3 (in Apendix A.2 along with the proof of the Theorem).
Corollary 5.2 (Routing regret).
Let $S^{*}(c)$ be an oracle shortlist (e.g., the best union-of-$k$-clusters set under $p_{c}$), define the routing regret as $\epsilon(c)\coloneqq p_{c}(S^{*}(c))-p_{c}(S(c))\geq 0$, then
$A(c)\geq(1-\beta)\left(p_{c}(S^{*}(c))-\epsilon(c)-\textnormal{TV}(p_{c}(\cdot\mid S),q_{c}(\cdot\mid S))\right).$
Note that FR-Spec has a static regret $\epsilon_{Fr}(c)=p_{c}(S^{*}(c))-p_{c}(S_{FR})$ while the corresponding quantity $\epsilon_{Dyn}(c)=p_{c}(S^{*}(c))-p_{c}(S_{Dyn}(c))$ is dynamic for Dynaspec. In particular, for an accurate router, then $\epsilon_{Dyn}(c)$ would be smaller on hard contexts such as code and rare symbols.
Theorem 5.3 (Expected accepted length).
Consider a block proposal of length $b$, and let the contexts along the draft be $c_{1},c_{2},\ldots,c_{b}$ such that each depends on previously proposed tokens. Let $A_{i}\coloneqq A(c_{i})$ be the overlap at step $i$. Let $L\in\{0,1,\ldots,b\}$ be the number of consecutively accepted draft tokens before the first rejection. If each step uses a proposal $\tilde{q}_{c_{i}}$ and the acceptance test is the standard speculative accept/reject, so that the probability of accepting token $i$ given it was reached is atleast $A_{i}$, then
$\mathbb{E}[L]\geq\sum_{i=1}^{b}\prod_{j=1}^{i}A_{j}.$ (6)
The proof is given appendix A.2. Plugging in $A_{j}=A_{\text{FR}}(c_{j})$ or $A_{\text{Dyn}}(c_{j})$ gives a method to compare the mean accepted lengths. The multiplicative form for the expected accepted length as highlighted in Theorem 5.3 shows that by having a larger pool of tokens to choose from, in the earlier steps of the drafting process, Dynaspec can outperform on the contexts where static $S_{Fr}$ misses mass. The above analysis also provides an analytical justification to the choice of having a position aware budgeting strategy (Equation equation 3) used in the router design.
@@ 6.1 Experimental Settings
We adhere to the experimental setup and evaluation of the Eagle (Li et al., 2024b, a) family and FR-Spec (Zhao et al., 2025). We implement the DynaSpec in the FR-Spec  [^1] https://github.com/thunlp/FR-Spec framework. All components are implemented in native C++ and CUDA with FR-Spec as the baseline, since Python call overhead would otherwise negate the latency gains from draft-time acceleration.
Datasets and Models. We benchmark DynaSpec over a diverse mix of seven tasks. Spec-Bench (Xia et al., 2024) supplies six of them—Machine Translation (WMT14 DE to EN) (Bojar et al., 2014), Multi-turn Conversation (MT-Bench) (Zheng et al., 2023), RAG and QA from Natural Questions (Kwiatkowski et al., 2019), Math from GSM8K (Cobbe et al., 2021), and Summarization (CNN/DailyMail) (Nallapati et al., 2016)—each evaluated on 80 prompts. The seventh task is Code, for which we employ HumanEval (Chen et al., 2021) with 164 problems. We bound generations to $1024$ tokens for Spec-Bench subtasks and 512 tokens for HumanEval. We used Llama series of models, specifically Llama-33-8B-Instruct, which uses a vocabulary size of $128K$ for our experiments and Qwen-2-7B.
Evaluation metrics and Hardware Settings. We evaluate using metrics: #Mean Acceptance Length, the average number of tokens committed per draft–verification cycle. #Mean Acceptance Length is hardware-agnostic but does not capture the overhead of drafting. Our primary baseline is FR-Spec (Zhao et al., 2025), a static subset–selection approach built on Eagle-2 (Li et al., 2024a); in contrast, DynaSpec performs dynamic subset selection atop Eagle-2 (Li et al., 2024a). All experiments were run on a single NVIDIA A6000 GPU.
Table 1: Mean accepted tokens for Llama-3-8B and Qwen-2-7B across seven datasets. vocab size indicates the shortlisted number of tokens. Full Vocab corresponds to the EAGLE-2 baseline; FR-Spec uses 32k. For DynaSpec, shortlist size varies by dataset.
[…]
Figure 3: Mean accepted tokens comparison for fixed top-$k$ clusters represented by DynaSpec-F vs. Position-aware top-$k$ clusters represented by DynaSpec-PA across seven datasets using Llama-3-8B model. The mean vocabulary sizes for DynaSpec-PA and DynaSpec-F are 28.7k and 29.2k, respectively.
@@
6.2 Mean Acceptance Length
We evaluate mean accepted length–the average number of draft tokens committed per draft verification cycle– across seven benchmarks. Table 1 reports results for Llama-3-8B-Instruct and Qwen-2-7B under three settings: (i) full-vocabulary Eagle-2 (upper bound), (ii) FR-Spec with a fixed 32K shortlist, and (iii) DynaSpec with a context-dependent shortlist whose average size is smaller than 32K.
Llama-3-8B. DynaSpec improves the average accepted length from 3.64 $\rightarrow$ 3.83. tokens/step (+5.2%) while using a smaller mean shortlist ($\sim 28K$ vs FR-Spec’s fixed 32K). Relative to full-vocabulary decoding (3.89), DynaSpec recovers $\sim 98.4\%$ of the upper bound, whereas FR-Spec recovers 93.6%. Gains are consistent across tasks, while Code setting shows the largest gap: static frequency truncation drops substantially below full-vocab (3.92), while DynaSpec nearly matches it (3.86), indicating that context-aware routing better preserves probability mass on rare identifiers, symbols, and formatting tokens that are systematically under-represented in frequency-ranked lists.
Qwen-2-7B. We observe the same pattern: DynaSpec improves the average accepted length from 3.37 $\rightarrow$ 3.53. while using an even smaller average shortlist ($\sim$ 20K vs 32K). Compared to full-vocabulary decoding (3.59), this corresponds to $\sim$ 98.3% recovery for DynaSpec vs. $\sim$93.9% for FR-Spec. Per-task improvements mirror the Llama results (notably Conversation 3.52 $\rightarrow$ 3.76 and Code 3.52 $\rightarrow$ 4.08, suggesting the benefit of dynamic, context-conditioned supports is robust across model families and vocabulary sizes.
Overall, these results support the central claim: dynamic, context-aware shortlisting can match (or closely approach) full-vocabulary acceptance while operating at substantially smaller effective vocabularies than fixed shortlists, which is exactly the regime needed to translate draft-head savings into end-to-end speculative decoding speedups, as described in the next section.
6.3 Decoding Speed
We measure end-to-end throughput (tokens/sec) for Llama-3-8B-Instruct on a single NVIDIA A6000, using a CUDA/C++ implementation to avoid Python overhead. We compare Target-only decoding, full-vocabulary Eagle-2, FR-Spec-32k (static shortlist), and DynaSpec-27k (dynamic routed shortlist).
[…]
6.4 Position-Aware Vocabulary Shortlisting
For the position-aware strategy, the results are presented in Figure 3. Using a position-aware subset size outperforms a fixed size across all draft steps, as the first few tokens have a disproportionate impact on acceptance and throughput, an observation consistent with prior findings in Gumiho Li et al. (2025). Allocating a larger cluster budget to early tokens preserves acceptance where it matters most, while decaying the shortlist later reduces indexed-GEMM cost without affecting verification. As shown in Figure 3, this adaptive schedule yields higher mean accepted lengths at comparable or smaller shortlist sizes. Table 4 (in Appendix) further reports results for combining the position-aware schedule with FR-Spec, which likewise improves mean acceptance and reduces average shortlist size. Overall, these results confirm that position awareness complements context-aware routing by focusing computation on the most influential draft steps and amortizing overhead in later ones.
7 Conclusion
We address a central bottleneck in speculative decoding for large-vocabulary LLMs—the drafter’s $\mathcal{O}(|V|d)$ output head. DynaSpec replaces static shortlists with a context-aware dynamic head that routes each prefix to a small number of token clusters and evaluates only over their union, while verification remains exact on the full vocabulary. Our analysis shows that context-conditioned supports strictly dominate any fixed subset in expected acceptance, and BCE-with-top-$k$ routing maximizes a lower bound on retained target mass. Systems-wise, we fuse index selection with the head GEMM and run routing in parallel, reducing draft-time cost. Empirically, DynaSpec consistently improves speculative decoding efficiency across standard benchmarks. It also achieves sustains high acceptance quality with substantially smaller candidate sets, while remaining plug-in compatible with EAGLE-style pipelines.
Impact Statement
This work improves the computational efficiency of LLM inference by reducing the drafter-side cost in speculative decoding through context-aware dynamic shortlisting, while leaving the target model’s full-vocabulary verification unchanged and therefore not altering the correctness of the decoding procedure.
Positive impacts include lower latency and potentially reduced energy/compute cost per generated token for deployments that already use speculative decoding. At the same time, efficiency improvements can also lower the marginal cost of generating large volumes of text, which may amplify downstream misuse (e.g., spam or automated content generation) depending on the application and surrounding safeguards. DynaSpec itself is a systems/decoding contribution and does not add new capabilities beyond faster generation; responsible deployment should rely on standard policy, access control, and content-moderation measures at the application level.
Appendix A Appendix
A.1 Dynaspec algorithmic procedure